OpenStudy (anonymous):

At 25 °C, what is the hydroxide ion concentration, [OH–], in an aqueous solution with a hydrogen ion concentration of [H ] = 1.9 × 10–4 M?

OpenStudy (surry99):

What is the relationship between H+, OH- and Kw in a solution?

OpenStudy (anonymous):

Kw = [H+] x [OH-] = 1.0 x 10-14

OpenStudy (jfraser):

so rearrange that equation to solve for [OH-] and plug in the values you know

OpenStudy (jfraser):

\[K_W = [H^{+1}]*[OH^{-1}] = 1*10^{-14}\]then\[[OH^{-1}] = \frac{K_W}{[H^{+1}]}\]

OpenStudy (anonymous):

so I got 2.7x10-15 is that correct?

OpenStudy (jfraser):

the hydroxide concentration must be larger than the value of KW. I suspect you made an error in your scientific notation in your calculator

OpenStudy (anonymous):

so, I divided 1x10-14 by 3.7 is that wrong? I got 3.7 by -log(1.9 x 10-4) =3.7

OpenStudy (jfraser):

you combined functions, you should have kept the [H+] as it was, \(1.9*10^{-4}\)

OpenStudy (jfraser):

so that \[[OH^{-1}] = \frac{1*10^{-14}}{[1.9*10^{-4}]}\]

OpenStudy (anonymous):

ok so 1x10-14/1.9x10-4=5.3x10-19

OpenStudy (surry99):

Try your calculation again there is a math mistake

OpenStudy (jfraser):

your order of operations isn't right.

OpenStudy (jfraser):

are you using a cell phone calculator or a real one?

OpenStudy (anonymous):

it should just be a simple cross multiplication problem right? 10.3/1=1x10-14/1.9x10-4 if so then this would = 5.5x10-18?

OpenStudy (surry99):

you should be solving the problem as indicated by @jfraser: [OH−1]=1∗10−14/[1.9∗10−4]

OpenStudy (anonymous):

i'm sorry i thought i was. oh=1x10-14/1.9x10-4 = 5.3x10-19 but i was told that was wrong. so then i put in a value for oh and that too is wrong.... what am i missing???? lastly in your response all that came through were a bunch of ????

OpenStudy (surry99):

[OH−1]=[1∗10−14]/[1.9∗10−4] = 5.26*10^-11

OpenStudy (anonymous):

i used a different calculator and I get 5.26x10-19 hmmm

OpenStudy (jfraser):

the way you're plugging in values must be wrong. When you put in values, how do you tell the calculator that it's scientific notation?

OpenStudy (anonymous):

actually i figured it out. I kept 1.9x10-4 it should have been 1.9x104